November 1996  ICTCM IX (Reno)  4

Example: The Parachute Problem

Abell and Braselton (Saunders, 1996, p. 140)

A parachutist weighing 60 kg falls from a plane (that is, v_{0} = 0) and is subjected to air resistance equivalent to F_{R} = 10v. After one minute, the parachute is opened so that the parachutist is subjected to air resistance equivalent to F_{R} = 100v.

1. What is the parachutist’s velocity when the parachute is opened?

2. What is the parachutist’s velocity v(t) after the parachute is opened?

3. What is the parachutist’s limiting velocity? How does this compare with the limiting velocity if the parachute does not open?